OpenStudy (anonymous):

Find all values of θ within the interval 0<θ

OpenStudy (marissalovescats):

Welcome to OpenStudy @Questionmachine! :)

OpenStudy (anonymous):

The equation can be expanded to:\[ \frac{\sin^2\theta}{\cos^2\theta} +\frac{\cos^2\theta}{\sin^2\theta} = 2 \]

OpenStudy (anonymous):

Multiply both sides by \(\sin^2\theta\cos^2\theta\) and get \[ \sin^4\theta+\cos^4\theta=2\sin^2\theta\cos^2\theta \]Interestingly enough... this is similar to \[ (a-b)^2 = a^2-2ab+b^2 \]

OpenStudy (anonymous):

So it simplifies to: \[ \sin^2\theta - \cos^2\theta = 0 \implies \sin^2\theta = \cos^2\theta \]

OpenStudy (anonymous):

Or you can use double angle formula \[ \sin^2\theta - \cos^2\theta =0 \implies (1- \cos^2\theta) - \cos^2\theta =0 \implies 2\cos^2\theta = 1 \]

OpenStudy (anonymous):

Thanks, Wio. But what is the final answer supposed to be ? What are the values that satisfy it ?

OpenStudy (anonymous):

Well, you get \[ |\cos\theta| = \frac{1}{\sqrt2} \]So you have to find values where \(0\lt \theta \lt \pi\)

OpenStudy (anonymous):



OpenStudy (anonymous):

So, the answer is 45 ? Sorry. I am a tad too slow at catching stuff.